HIGH-RESOLUTION INFRARED SPECTROSCOPY OF BUTADIENE: TRANS AND GAUCHE RO-TAMERS.

\underline{NORMAN C. CRAIG}, Department of Chemistry, Oberlin College, Oberlin, OH 44074; ROBERT L. SAMS, ESML, Pacific Northwest National Laboratory, Richland, WA 99352.

In a search for rotational structure of the gauche rotamer of butadiene, high-resolution (0.002 cm^{−1}) infrared spectra were recorded at a long path length in regions that grew in intensity when the temperature was raised from 0^{∘}C to 50^{∘}C. These regions were also ones where Huber-Wälchli and Günthard had observed bands for a second, higher energy rotamer, when hot butadiene was frozen in an argon matrix at 4 K.^{a} From medium-resolution infrared spectra, DeMaré and coworkers had concluded that a band at 749 cm^{−1} was from the gauche rotamer.^{b} Our high-resolution analysis of this largely A-type band indicates it is, however, probably a difference tone [ν_{15}(b_{g}) - ν_{13}(a_{u})] of the trans rotamer. Attempts to identify other subband series attributable to the gauche rotamer, which is present to only about 5% at room temperature,^{c} failed. Much of the searching occurred in the wings of the 525-cm^{−1} band for ν_{12}(a_{u}) of the trans rotamer, which has been analyzed for the first time. In addition to rotational transitions associated with the ν_{12} fundamental, the rotational structure in hot bands was also analyzed.

aP. Huber-Wälchli, Hs. H. Günthard, Spectrochim. Acta 1981, 37A, 285.

bG. R. DeMaré, Yu. N. Panchenko, J. Vander Auwera, J. Phys. Chem. A 1997, 101 , 3998.

cJ. Saltiel, D. F. Sears, Jr., A. M. Turek, J. Phys. Chem. A 2001, 105 , 7569.